What is the axis of symmetry of x^2+9x+4?
Answer by kiru_khandelwal(79)

If the equation of the parabola is of the form ax^2 + bx + c = y
then the axis of symmetry is x=-b/2a
In our case
x^2 + 9x + 4
=> a =1 b = 9 and c = 4
so axis of symmetry is x= -b/2a = -9/2*1 = -9/2
s = -9/2